THE ARE 20 STUDENTS AND THEY NEED TO BE ALLOCATED INTO 3 CLASSROOMS.IN HOW MANY WAYS CAN THIS BE DONE
1.ASSUMING THAT THE STUDENTS ARE IDENTICAL
2.ASSUMING THAT THE STUDENTS ARE DISTINGUISHABLE

Answer by khwang(438)

1.ASSUMING THAT THE STUDENTS ARE IDENTICAL
2.ASSUMING THAT THE STUDENTS ARE DISTINGUISHABLE
1. Like the possible solutions of the integral equations
x + y + z = 20 (z,y & z as three classrooms)
Arrange r = 20 possble combinations among n=3 elements.
There are H(n, r) = C(n+r-1, r) = C(22,20) = C(22,2)
=22*21/2 = 21*11 = 231
2. Each student has 3 choices, totally, there 3^20 ways.

Kenny